OpenStudy (anonymous):

can't remember the exact question, but I was given a number: 17^(1/4), and told to find the linear proximation.

OpenStudy (turingtest):

well what number close to 17 do we know the 4th root of exactly?

OpenStudy (anonymous):

16^1/4

OpenStudy (anonymous):

=2

OpenStudy (turingtest):

good, so what function will we be approximating?

OpenStudy (anonymous):

thats the thing, there was no function! It was part b of a question on my exam, but part A was totally unrelated to it... D:

OpenStudy (turingtest):

the idea of a linear approximation is that we create a function that we know a known value of to estimate a value close to it

OpenStudy (turingtest):

in this case we are asked to find the 4th root of a number we know the 4th root of a number close to it hence out function should be...?

OpenStudy (anonymous):

exactly, I asked my prof because I thought there was a typo, but she smiled and said I was on the right track..

OpenStudy (anonymous):

I don't understand..

OpenStudy (turingtest):

we are approximating 4th roots, does it not make sense to make use of the function\[f(x)=x^{1/4}\]???

OpenStudy (anonymous):

so, find the linear app. of that equation with the point given?

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

so what x value is our approximation based around?

OpenStudy (anonymous):

would that be the number given in the question??

OpenStudy (turingtest):

no, we need to base our estimate around an x value \(x_0\) for which we already know f(x) what is the x value that we already said we know f(x) of ?

OpenStudy (anonymous):

16

OpenStudy (turingtest):

yes, so \(f(x)=x^{1/4}\) and \(x_0=16\) for our purposes now we just need to fill in the pieces to get our linear approximation function...

OpenStudy (turingtest):

for \(x\approx x_0\) we have the linear approximation\[f(x)\approx f(x_0)+f'(x_0)(x-x_0)\]

OpenStudy (turingtest):

so we need all the pieces of the formula... what is \(f(x_0)\) ?

OpenStudy (anonymous):

2

OpenStudy (turingtest):

good what is \(f'(x)\) ?

OpenStudy (anonymous):

1/4(x^(-3/4))

OpenStudy (turingtest):

good, so what is \(f'(x_0)\) ?

OpenStudy (anonymous):

1/32?

OpenStudy (turingtest):

excellent, not plug into the formula\[f(x)\approx f(x_0)+f'(x_0)(x-x_0)\]what do you get?

OpenStudy (turingtest):

now*

OpenStudy (anonymous):

x/32 + 3/2?

OpenStudy (turingtest):

yep :) now we can estimate f(17), which comes out to be...?

OpenStudy (anonymous):

O.O you mean just plug it into the equation?! And get a 2.03??

OpenStudy (turingtest):

yep, that's what we're doing :)

OpenStudy (anonymous):

O.O did we just answer the question?!

OpenStudy (anonymous):

I mean you did.. and me ish

OpenStudy (turingtest):

yes you did; you really did all the work yourself, I just guided you notice the logic just so it doesn't seem magic...

OpenStudy (anonymous):

:D awesome thanks!

OpenStudy (anonymous):

which means I got that question wrong on my test D:

OpenStudy (turingtest):

for \(x\approx x_0\)\[f(x)\approx f(x_0)+f'(x_0)(x-x_0)\]in our case we got\[f(x)\approx\frac x{32}+\frac32\]so\[f(17)\approx\frac {17}{32}+\frac32=\frac{65}{32}\] since \(f(x)=x^{1/4}\) this statement says\[17^{1/4}\approx\frac{65}{32}\]which is what we wanted to know

OpenStudy (turingtest):

the linear approximation gives 2.03125 and my calculator gives 2.03054... so they are pretty darn close sorry to hear about your test, better luck next time!

OpenStudy (anonymous):

D: thanks anyways! btw (y) cowboy bebop!

OpenStudy (turingtest):

yeah, it's a classic! Spike and I have the same last name too :)

OpenStudy (anonymous):

O.O! sorry, had to check your profile, thats so cool

OpenStudy (turingtest):

:)